# Double Robustness of Local Projections

### v1

@@ Abstract:
We consider impulse response inference in a locally misspecified stationary vector autoregression (VAR) model. The conventional local projection (LP) confidence interval has correct coverage even when the misspecification is so large that it can be detected with probability approaching 1. This follows from a “double robustness” property analogous to that of modern estimators for partially linear regressions. In contrast, VAR confidence intervals dramatically undercover even for misspecification so small that it is difficult to detect statistically and cannot be ruled out based on economic theory. This is because of a “no free lunch” result for VARs: the worst-case bias and coverage distortion are small if, and only if, the variance is close to that of LP. While VAR coverage can be restored by using a bias-aware critical value or a large lag length, the resulting confidence interval tends to be at least as wide as the LP interval.
@@ 1 Introduction
While this sentiment has been echoed in influential reviews (e.g., see Ramey, 2016; Nakamura and Steinsson, 2018; Jordà, 2023), there so far exist essentially no theoretical results on the relative robustness of LP and VAR inference procedures to misspecification. Plagborg-Møller and Wolf (2021) and Xu (2023) show that the two estimators are in fact asymptotically equivalent—and thus equally robust to misspecification—in a general VAR($\infty$) model if the estimation lag length diverges to infinity with the sample size. However, this result does not directly speak to the empirically relevant case where researchers employ small-to-moderate lag lengths to preserve degrees of freedom. Applied researchers must therefore base their choice of inference procedure on empirically calibrated simulation studies (Kilian and Kim, 2011; Li, Plagborg-Møller, and Wolf, 2024).
@@
The main contribution of this paper is to provide a formal proof of the claim in Jordà (2005) that conventional LP confidence intervals for impulse responses are surprisingly robust to misspecification. At the same time, we show that VAR confidence intervals are unreliable in the face of even small amounts of misspecification.
@@
To formalize our results, we consider a large class of stationary data generating processes (DGPs) that are well approximated by a finite-order SVAR model, but subject to local misspecification in the form of an asymptotically vanishing moving average (MA) process, of potentially infinite order. This class of models covers many types of dynamic misspecification, such as under-specification of the lag length, failure to include relevant control variables, inappropriate aggregation, measurement error, and omitted nonlinearities. In particular, such an environment is consistent with essentially all linearized structural macroeconomic models (Fernández-Villaverde, Rubio-Ramírez, Sargent, and Watson, 2007). Intuitively, our set-up with local misspecification captures the idea that finite-order VAR models provide a good—but not perfect—approximation of reality.
@@
In this setting, we prove that the conventional LP confidence interval has correct (pointwise) asymptotic coverage even for local misspecification that is of such a large magnitude that it can be detected with probability 1 in large samples. This robustness property requires that we control for those lags of the data that are strong predictors of the outcome or impulse variables, but crucially for applied work, the omission of lags with small-to-moderate predictive power does not threaten coverage. We argue that our result can be interpreted as a consequence of the double robustness of the LP estimator, which is analogous to the double robustness of modern partially linear regression estimators in the literature on debiased machine learning (Newey, 1990; Chernozhukov, Chetverikov, Demirer, Duflo, Hansen, Newey, and Robins, 2018; Chernozhukov, Escanciano, Ichimura, Newey, and Robins, 2022).
@@
In stark contrast to LP, small amounts of misspecification cause conventional VAR confidence intervals for impulse responses to suffer from severe under-coverage asymptotically. We derive analytically the worst-case bias and coverage of VARs over all possible misspecification processes, subject to a constraint only on the overall magnitude of the misspecification. A “no free lunch” result for VARs emerges: the worst-case bias and coverage distortion are small if, and only if, the asymptotic variance is close to that of LP. This worst-case result is practically relevant—VAR confidence intervals severely undercover even when the misspecification term: (i) is small in magnitude; (ii) has dynamic properties that cannot be ruled out ex ante based on economic theory; and (iii) is difficult to detect ex post with model specification tests. While the coverage can be restored by using a larger bias-aware critical value (Armstrong and Kolesár, 2021) or by using a large number of lags for estimation, the resulting confidence intervals are so wide that one may as well report the LP interval.
@@
We demonstrate through simulations that our asymptotic results are useful to understand the finite-sample trade-off between LP and VAR confidence intervals. We first consider a univariate ARMA(1,1) model. While LP confidence intervals here have coverage close to the nominal level, confidence intervals based on short-lag autoregressions severely undercover when the MA coefficient is only moderately large. In fact, in this simple case the implied MA misspecification is very close to the theoretical worst case. We next run simulations based on the medium-scale structural macroeconomic model of Smets and Wouters (2007). The researcher estimates the impulse response function of inflation with respect to an observed cost-push shock, with lag length selected by the Akaike Information Criterion (AIC). VAR confidence intervals materially undercover—particularly at medium and long horizons—while LP again attains close to nominal coverage. Consistent with our theoretical results, increasing the estimation lag length ameliorates the VAR coverage, but at the cost of delivering confidence intervals as wide as those of LP. [^1] We obtain similar results for simulations based on other observed shocks (including monetary shocks) and other shock identification schemes (including recursive identification).
Other literature.
@@
Relative to the previously cited simulation studies of LPs and VARs, we derive analytical results on the worst-case asymptotic properties of these inference procedures that hold for a wide range of stationary, locally misspecified VAR models. The empirically calibrated simulations in Li, Plagborg-Møller, and Wolf (2024) suggest a stark bias-variance trade-off between LP (low bias, high variance) and VAR estimators (moderate bias, low variance). The reason behind the theoretical superiority of LP proved in this paper is that, if the objective is to construct confidence intervals, then even a moderate amount of VAR bias cannot be tolerated, as it causes the VAR confidence interval to be poorly centered. Effectively, a concern for confidence interval coverage induces a large weight on bias in the researcher’s objective function, justifying the use of LP despite its higher variance.
@@
The robustness of LP to misspecification discussed here—with stationary data and at fixed horizons—is conceptually and theoretically distinct from the robustness of LP to the persistence in the data and the length of the impulse response horizon shown by Montiel Olea and Plagborg-Møller (2021). Nevertheless, it turns out that controlling for lags (“lag augmentation”) is key to all the robustness properties established in Montiel Olea and Plagborg-Møller (2021) and the present paper.
@@
We also build upon previous research into misspecified VAR models, uncovering novel results about the robustness of LPs and the worst-case properties of VAR procedures. Braun and Mittnik (1993) derive expressions for the probability limits of VAR estimators under global MA misspecification, but do not provide analytical results on the performance of inference procedures. Schorfheide (2005) characterizes the asymptotic mean squared errors of iterated and direct multi-step forecasts in a reduced-form VAR model with MA terms of order $T^{-1/2}$. Müller and Stock (2011) construct Bayesian forecast intervals in a locally misspecified univariate AR model. Relative to these papers, we contribute by: (i) focusing on structural impulse responses rather than forecasting; (ii) allowing for more general rates of local misspecification, key to uncovering the double robustness of LP; and (iii) deriving simple analytical formulas for the worst-case bias and coverage of VARs. As such, our results formalize concerns by some applied practitioners about the lack of VAR robustness (e.g., Chari, Kehoe, and McGrattan, 2008; Nakamura and Steinsson, 2018).
@@
Whereas our paper deals with bias imparted by dynamic misspecification, our asymptotics abstract from the order-$T^{-1}$ bias of LPs and VARs that is due to persistence in the data (Pope, 1990; Kilian, 1998; Herbst and Johannsen, 2023).
@@
Section 2 provides an overview of our main results in a univariate model. Section 3 introduces the general local-to-SVAR($p$) model, defines LP and VAR estimators, and proves the robustness of LP and fragility of VAR confidence intervals. Section 4 derives analytically the worst-case bias and coverage of VARs, and shows that bias-aware VAR confidence intervals tend to be wider than the LP interval. Section 5 demonstrates the practical relevance of our results through simulations. Section 6 concludes. Proofs and supplementary results are relegated to the appendix and an online supplement.
@@
2 Overview of results
In this section we illustrate our main results in a simple special case, namely a univariate autoregressive (AR) model subject to misspecification in the form of a small moving average (MA) process. In the subsequent sections we will show that all qualitative results—as well as many quantitative ones—from the simple model carry over to a general class of multivariate SVAR models with local MA misspecification.
Model and assumptions.
Suppose the data $\{y_{t}\}_{t=1}^{T}$ satisfies a univariate, stationary local-to-AR(1) model222$\{y_{t}\}$ is a triangular array, but we suppress the dependence on $T$ in the notation for convenience.
$y_{t}=\rho y_{t-1}+[1+T^{-\zeta}\alpha(L)]\varepsilon_{t}\quad\text{for all }t,$ (2.1)
where $\alpha(L)=\sum_{\ell=1}^{\infty}\alpha_{\ell}L^{\ell}$ is a potentially infinite-order lag polynomial. This model captures the idea that the time series dynamics of the data are well approximated by a simple autoregressive model (here an AR(1)) driven by an unobserved white noise shock $\varepsilon_{t}$, but with a small amount of misspecification in the form of an MA process $T^{-\zeta}\alpha(L)\varepsilon_{t}$. The misspecification is asymptotically small in the sense that the MA coefficients converge to zero at the rate $T^{-\zeta}$, though the misspecification may still affect the properties of estimators, as shown by Schorfheide (2005) and as demonstrated below. We argue below in Section 3.1 that MA misspecification of this form can capture many empirically relevant types of dynamic misspecification, such as under-specification of the lag length, missing control variables, inappropriate aggregation, and measurement error. We consider local rather than global misspecification in the spirit of local power analysis (e.g., Rothenberg, 1984), since this makes the bias-variance trade-off between the AR and LP estimators matter even asymptotically as the sample size diverges, allowing us to make tractable analytical comparisons between these two procedures.
Assumption 2.1.
For each $T$, $\{y_{t}\}_{t\in\mathbb{Z}}$ is the stationary solution to the equation (2.1), where the shocks and parameters satisfy the following conditions:
  1. i)
    $\varepsilon_{t}\overset{i.i.d.}{\sim}(0,\sigma^{2})$ with $\sigma^{2}>0$ and $E(\varepsilon_{t}^{4})<\infty$.
  2. ii)
    $\rho\in(-1,1)$.
  3. iii)
    $\alpha(L)$ is absolutely summable.
  4. iv)
    $\zeta>1/4$.
Absolute summability of $\alpha(L)$ is a weak regularity condition ensuring the MA($\infty$) process $\alpha(L)\varepsilon_{t}$ is well-defined (Brockwell and Davis, 1991, Proposition 3.1.1). The significance of the assumption that misspecification vanishes faster than $T^{-1/4}$ will become clear below.
The parameter of interest is the impulse response of $y_{t+h}$ with respect to $\varepsilon_{t}$ according to the model (2.1):
$\theta_{h,T}\equiv E[y_{t+h}\mid\varepsilon_{t}=1]-E[y_{t+h}\mid\varepsilon_{t}=0]=\rho^{h}+T^{-\zeta}\sum_{\ell=1}^{h}\rho^{h-\ell}\alpha_{\ell},\quad h=0,1,2,\dots.$
The first term is the usual AR(1) formula, while the second term arises from the MA component. Importantly, and consistent with our focus on the consequences of dynamic misspecification, we do not treat the VAR misspecification as non-classical measurement error that should be ignored for structural analysis; instead, the true causal model has an ARMA form (with small but potentially non-zero MA terms), and we care about the full transmission mechanism of shocks in this model.
Estimators.
We will consider two estimators of the impulse response of interest $\theta_{h,T}$ using the data $\{y_{t}\}_{t=1}^{T}$:
  1. 1.
    The LP estimator $\hat{\beta}_{h}$ is obtained from an OLS regression
    $y_{t+h}=\hat{\beta}_{h}y_{t}+\hat{\gamma}_{h}y_{t-1}+\hat{\xi}_{h,t},$ (2.2)
    where we control for one lag of the data, and $\hat{\xi}_{h,t}$ is the OLS residual.
  2. 2.
    The AR estimator $\hat{\delta}_{h}$ is given by
    $\hat{\delta}_{h}=\hat{\rho}^{h},\quad\text{where}\quad\hat{\rho}=\frac{\sum_{t=2}^{T}y_{t}y_{t-1}}{\sum_{t=2}^{T}y_{t-1}^{2}}.$
The two estimators coincide at the impact horizon, $\hat{\beta}_{0}=\hat{\delta}_{0}$. Conventional confidence intervals based on both these estimators would have correct asymptotic coverage in a well-specified AR(1) model. However, the presence of the MA term in the model (2.1) means that, in principle, both the LP and VAR estimators ought to control for infinitely many lags of the data rather than just one. Nevertheless, we shall now see that this dynamic misspecification has much more serious consequences for the AR procedure than for LP.
Robustness of LP.
Our first main result is that the limiting distribution of the LP estimator is remarkably robust to the presence of local misspecification.
Proposition 2.1.
Under Assumption 2.1,
$\hat{\beta}_{h}-\theta_{h,T}=\frac{1}{\sigma^{2}}\frac{1}{T}\sum_{t=1}^{T}\xi_{h,t}\varepsilon_{t}+o_{p}(T^{-1/2}),$
where $\xi_{h,t}\equiv\sum_{\ell=1}^{h}\rho^{h-\ell}\varepsilon_{t+\ell}$ is the $h$-step-ahead forecast error.
Since the above asymptotic representation is invariant to the MA misspecification, the limiting distribution of LP is the same as it would be in a correctly specified AR(1) model (provided $\zeta>1/4$, as imposed in Assumption 2.1). Though this robustness of LP is with respect to local (i.e., asymptotically vanishing) misspecification, it is still quantitatively meaningful, as MA terms of order $T^{-\zeta}$ with $\zeta\in(1/4,1/2)$ can be detected with probability 1 asymptotically by conventional AR model specification tests, such as the Hausman test considered in Section 3.3 below.
Why is LP robust to misspecification of such large magnitude? We will offer two mathematically equivalent pieces of intuition, with our discussion here deliberately heuristic. First, the classic omitted variable bias formula suggests that the bias of the LP impulse response estimator $\hat{\beta}_{h}$ in the regression (2.2) is proportional to the product of two factors: (i) the direct effect of omitted lags on $y_{t+h}$, and (ii) the covariance of the residualized regressor of interest $y_{t}-E[y_{t}\mid y_{t-1}]=\varepsilon_{t}+O_{p}(T^{-\zeta})$ with the omitted lags. The factor (i) is of order $T^{-\zeta}$ in the local-to-AR(1) model (2.1). The factor (ii) is also of order $T^{-\zeta}$, since $\varepsilon_{t}$ is uncorrelated with any lagged data. Hence, the OVB is of order $T^{-2\zeta}=o(T^{-1/2})$ under the assumption $\zeta>1/4$, and thus the bias of the estimator is negligible relative to the standard deviation, which is of order $T^{-1/2}$ as in the correctly specified case.
The preceding intuition is a special case of the double robustness property of partially linear regressions (see Example 1 in Chernozhukov, Escanciano, Ichimura, Newey, and Robins, 2022), which we will now argue applies also to LP, again settling for a heuristic argument. Consider any dynamic model (for example a VARMA($p,q$)) that implies the following local projection representation:
[…]
Here $\theta_{0,h}$ is the true impulse response, and $\gamma_{0}(\cdot)$ is a function of lagged data. Define $\nu_{0}(y^{t-1})\equiv E[y_{t}\mid y^{t-1}]$. By applying the Frisch-Waugh lemma to the regression (2.2), the LP estimator $\hat{\beta}_{h}$ is the sample analogue of the solution $\theta_{0,h}$ to the moment condition
$E[\{y_{t+h}-\theta_{0,h}y_{t}-\gamma_{0}(y^{t-1})\}\{y_{t}-\nu_{0}(y^{t-1})\}]=0.$
If we evaluate the moment on the left-hand side at arbitrary functions $\gamma(\cdot)$ and $\nu(\cdot)$ rather than at the true ones $\gamma_{0}(\cdot)$ and $\nu_{0}(\cdot)$, a simple calculation shows that it equals $E[\{\gamma_{0}(y^{t-1})-\gamma(y^{t-1})\}\{\nu_{0}(y^{t-1})-\nu(y^{t-1})\}]$. [^3] We can write the moment as $E[\{y_{t+h}-\theta_{0,h}y_{t}-\gamma_{0}(y^{t-1})+\gamma_{0}(y^{t-1})-\gamma(y^{t-1})\}\{y_{t}-\nu(y^{t-1})\}]=E[\{\gamma_{0}(y^{t-1})-\gamma(y^{t-1})\}\{y_{t}-\nu_{0}(y^{t-1})+\nu_{0}(y^{t-1})-\nu(y^{t-1})\}]$, since $y_{t+h}-\theta_{0,h}y_{t}-\gamma_{0}(y^{t-1})=\xi_{h,t}$ is independent of $y^{t}$. The claim now follows from $E[y_{t}-\nu_{0}(y^{t-1})\mid y^{t-1}]=0$ by definition of $\nu_{0}(\cdot)$. Hence, the moment condition is satisfied at the true impulse response parameter $\theta_{0,h}$ as long as either $\gamma=\gamma_{0}$ or $\nu=\nu_{0}$, making the LP estimator doubly robust: it is consistent if we correctly specify either the controls $\gamma(y^{t-1})$ in the outcome equation or the controls $\nu(y^{t-1})$ in the implicit first-stage regression that isolates the shock $\varepsilon_{t}=y_{t}-\nu(y^{t-1})$. Whereas in the present univariate model the choices of $\gamma(\cdot)$ and $\nu(\cdot)$ are necessarily linked, in the general multivariate model in Section 3 they are decoupled, and as a result it is possible to correctly specify the controls in the outcome equation while mis-identifying the shock, and vice versa. As a consequence of double robustness, and as argued more generally by Chernozhukov, Chetverikov, Demirer, Duflo, Hansen, Newey, and Robins (2018) (and confirmed by our proof), it turns out that estimation error in $\gamma_{0}$ and $\nu_{0}$ only affects the asymptotic distribution of the doubly robust estimator $\hat{\beta}_{h}$ through the product of the estimation errors $\|\hat{\gamma}-\gamma_{0}\|\times\|\hat{\nu}-\nu_{0}\|$. In our local-to-AR(1) model (2.1), both terms in this product are of order $T^{-\zeta}$ due to the omitted lags. The product is then of order $T^{-2\zeta}=o(T^{-1/2})$ and thus asymptotically negligible, consistent with our earlier intuition.
Fragility of AR.
Our second main result is that this (double) robustness of LP to misspecification stands in stark contrast to the fragility of the AR estimator.
Proposition 2.2.
Under Assumption 2.1,
$\hat{\delta}_{h}-\theta_{h,T}=T^{-\zeta}\operatorname{aBias}(\hat{\delta}_{h})+\frac{h\rho^{h-1}(1-\rho^{2})}{\sigma^{2}}\frac{1}{T}\sum_{t=1}^{T}\varepsilon_{t}\tilde{y}_{t-1}+o_{p}(T^{-\zeta}+T^{-1/2}),$ (2.3)
where $\tilde{y}_{t}=(1-\rho L)^{-1}\varepsilon_{t}$ satisfies a correctly specified AR(1) model with $\alpha(L)=0$, and the asymptotic bias is
$\operatorname{aBias}(\hat{\delta}_{h})\equiv\underbrace{h\rho^{h-1}}_{\frac{\partial(\rho^{h})}{\partial\rho}}\underbrace{(1-\rho^{2})\sum_{\ell=1}^{\infty}\rho^{\ell-1}\alpha_{\ell}}_{\operatorname{aBias}(\hat{\rho})=\frac{\operatorname*{Cov}(\alpha(L)\varepsilon_{t},\tilde{y}_{t-1})}{\operatorname*{Var}(\tilde{y}_{t-1})}}-\underbrace{\sum_{\ell=1}^{h}\rho^{h-\ell}\alpha_{\ell}}_{\theta_{h,T}-\rho^{h}}.$ (2.4)
The convergence rate of the AR estimator is $T^{-\min\{1/2,\zeta\}}$, potentially slower than the $T^{-1/2}$ rate achieved by LP. This is because the AR estimator suffers from an asymptotic bias of order $T^{-\zeta}$. This bias is due to two forces: first, the AR(1) coefficient $\hat{\rho}$ is biased due to the endogeneity caused by the MA terms, and second, the AR estimator extrapolates the $h$ impulse response based on a parametric formula $\hat{\rho}^{h}$ that does not hold exactly in the true model (2.1). The second term on the right-hand side of the asymptotic representation (2.3) is the same as it would be in the correctly specified AR(1) model (with $\alpha(L)=0$), and therefore of order $T^{-1/2}$. As a consequence, the AR bias is only asymptotically negligible if $\zeta>1/2$, a much smaller degree of robustness than shown above for LP. The case $\zeta=1/2$ is of particular interest, as then the bias and standard deviation are of the same asymptotic order (see also Schorfheide, 2005). MA terms of order $T^{-1/2}$ can be detected with asymptotic probability strictly between 0 and 1 by specification tests, as will be shown below in Section 3.3.
Implications for inference.
Consider the conventional level-$(1-a)$ confidence intervals based on the LP and AR estimators, respectively:
$\operatorname{CI}(\hat{\beta}_{h})\equiv\left[\hat{\beta}_{h}\pm z_{1-a/2}\sqrt{\operatorname{aVar}(\hat{\beta}_{h})/T}\right],\quad\operatorname{CI}(\hat{\delta}_{h})\equiv\left[\hat{\delta}_{h}\pm z_{1-a/2}\sqrt{\operatorname{aVar}(\hat{\delta}_{h})/T}\right],$ (2.5)
where $z_{1-a/2}\equiv\Phi^{-1}(1-a/2)$ is the normal critical value, $\Phi(\cdot)$ is the standard normal cumulative distribution function, and the asymptotic variances $\operatorname{aVar}(\hat{\beta}_{h})$ and $\operatorname{aVar}(\hat{\delta}_{h})$ correspond to the variances of the leading stochastic terms in Propositions 2.1 and 2.2, yielding the same expressions as in the correctly specified case. [^4] Under Assumption 2.1, the summands of the leading stochastic terms are serially uncorrelated, so the asymptotic variances depend on the simple variance rather than the long-run variance. Then $\operatorname{aVar}(\hat{\beta}_{h})\equiv\sigma^{-4}\operatorname*{Var}(\xi_{h,t}\varepsilon_{t})=(1-\rho^{2h})/(1-\rho^{2})$ and $\operatorname{aVar}(\hat{\delta}_{h})\equiv\sigma^{-4}[h\rho^{h-1}(1-\rho^{2})]^{2}\operatorname*{Var}(\varepsilon_{t}\tilde{y}_{t-1})=(h\rho^{h-1})^{2}(1-\rho^{2}).$ While we use the asymptotic variances in the definition of the confidence intervals for mathematical simplicity, the results would be the same if we used the conventional (estimated) standard error formulas that apply under correct specification, since the latter are consistent under Assumption 2.1.
Corollary 2.1.
Under Assumption 2.1, $\lim_{T\to\infty}P(\theta_{h,T}\in\operatorname{CI}(\hat{\beta}_{h}))=1-a$. If moreover $\operatorname{aVar}(\hat{\delta}_{h})>0$ and $\operatorname{aBias}(\hat{\delta}_{h})\neq 0$, then $\lim_{T\to\infty}P(\theta_{h,T}\in\operatorname{CI}(\hat{\delta}_{h}))=0$ for $\zeta\in(1/4,1/2)$, and $\lim_{T\to\infty}P(\theta_{h,T}\in\operatorname{CI}(\hat{\delta}_{h}))<1-a$ for $\zeta=1/2$.
The corollary shows that the LP confidence interval robustly controls coverage. In contrast, except in the knife-edge case where the asymptotic bias equals 0, the AR confidence interval has coverage converging to 0 for misspecification of order $T^{-\zeta}$ with $\zeta\in(1/4,1/2)$, while the asymptotic coverage is strictly below the nominal level $1-a$ for $\zeta=1/2$. This is because the AR interval has the same width as in the correctly specified model, but it is incorrectly centered due to the bias.
Worst-case bias and coverage.
While we have shown that the AR confidence interval can severely under-cover asymptotically, it is natural to ask whether such poor performance is likely in practice. To answer this question, we set $\zeta=1/2$, ensuring that the bias-variance trade-off is non-trivial; otherwise the bias either dominates or is negligible asymptotically. Next, we restrict the misspecification to be small, in the sense that the noise-to-signal ratio $\operatorname*{Var}(T^{-1/2}\alpha(L)\varepsilon_{t})/\operatorname*{Var}(\varepsilon_{t})$ in the local-to-AR(1) model (2.1) is bounded by $M^{2}/T$ for a given $M$. Under this constraint, the worst-case VAR bias is strikingly simple.
Proposition 2.3.
Under Assumption 2.1, $\zeta=1/2$, and $\operatorname{aVar}(\hat{\delta}_{h})>0$,
$\max_{\alpha(L)\colon\operatorname*{Var}(\alpha(L)\varepsilon_{t})/\sigma^{2}\leq M^{2}}\frac{|\operatorname{aBias}(\hat{\delta}_{h})|}{\sqrt{\operatorname{aVar}(\hat{\delta}_{h})}}=M\sqrt{\frac{\operatorname{aVar}(\hat{\beta}_{h})}{\operatorname{aVar}(\hat{\delta}_{h})}-1}.$ (2.6)
This formula implies that there is no free lunch for the AR estimator: whenever the relative standard error $\sqrt{\operatorname{aVar}(\hat{\delta}_{h})/\operatorname{aVar}(\hat{\beta}_{h})}$ of the AR and LP estimators is small, the AR estimator is vulnerable to misspecification in the sense of a large worst-case bias, unless we are certain a priori that the noise-to-signal ratio $M^{2}/T$ is extremely small. It turns out that the exact same worst-case bias formula applies to multivariate locally misspecified VAR models with multiple lags, as shown below in Section 4.1. Hence, while it is possible to ameliorate the worst-case bias of AR estimators by increasing the lag length, this can only happen at the expense of increasing the variance.
As an implication of the above result, we show analytically in Section 4 that the AR confidence interval has severely distorted worst-case coverage when the relative standard error of AR relative to LP lies in an empirically relevant range. These severe distortions are achieved by MA misspecification processes that (i) are small, (ii) cannot be ruled out ex ante using economic theory, and (iii) are difficult to detect ex post using model specification tests. While it is possible to construct a valid bias-aware AR confidence interval with a larger critical value that takes into account the worst-case bias (following the general recipe of Armstrong and Kolesár, 2021), this interval tends to be wider than the equally valid conventional LP confidence interval unless the noise-to-signal ratio is restricted to be extremely small. We conclude that AR inference is only reliable when using a lag length that is so large that the procedure becomes equivalent with LP.
3 Robust local projections, fragile VARs
We now state our general model and assumptions, define the general LP and VAR impulse response estimators, derive the asymptotic representations of these estimators, and draw conclusions about their relative robustness to misspecification. While LP is shown to be robust to large amounts of misspecification, the VAR estimator is fragile in that it suffers from generically non-negligible asymptotic bias.

### v2

@@ Abstract:
We consider impulse response inference in a locally misspecified vector autoregression (VAR) model. The conventional local projection (LP) confidence interval has correct coverage even when the misspecification is so large that it can be detected with probability approaching 1. This result follows from a “double robustness” property analogous to that of popular partially linear regression estimators. In contrast, the conventional VAR confidence interval with short-to-moderate lag length can severely undercover, even for misspecification that is small, economically plausible, and difficult to detect statistically. There is no free lunch: the VAR confidence interval has robust coverage only if the lag length is so large that the interval is as wide as the LP interval.
@@ 1 Introduction
While this sentiment has been echoed in influential reviews (e.g., Ramey, 2016; Nakamura and Steinsson, 2018; Jordà, 2023), there so far exist essentially no theoretical results on the relative robustness of LP and VAR inference procedures to misspecification. Plagborg-Møller and Wolf (2021) and Xu (2023) show that the two estimators are in fact asymptotically equivalent—and thus equally robust to misspecification—in a general VAR($\infty$) model if the estimation lag length diverges to infinity with the sample size. However, this result does not directly speak to the empirically relevant case where researchers employ small-to-moderate lag lengths to preserve degrees of freedom. Applied researchers must therefore base their choice of inference procedure on empirically calibrated simulation studies (Kilian and Kim, 2011; Li, Plagborg-Møller, and Wolf, 2024).
@@
In this paper we provide a formal proof of Jordà’s claim that conventional LP confidence intervals for impulse responses are surprisingly robust to misspecification. For VAR confidence intervals, we on the other hand show that there is no free lunch: they are robust if, and only if, they are as wide as LP intervals asymptotically, as is the case when they feature a large number of lags. If the confidence interval is shorter, then it is necessarily unreliable.
@@
We consider a large class of stationary data generating processes (DGPs) that are well approximated by a finite-order SVAR model, but subject to local misspecification in the form of an asymptotically vanishing moving average (MA) process, of potentially infinite order. This class is consistent with essentially all linearized structural macroeconomic models and covers many types of dynamic misspecification, such as under-specification of the lag length, failure to include relevant control variables, inappropriate aggregation, and measurement error. Intuitively, with this set-up we capture the idea that finite-order VAR models provide a good but imperfect approximation of reality.
@@
In this setting, we prove that the conventional LP confidence interval has correct (pointwise) asymptotic coverage even for local misspecification that is of such a large magnitude that it can be detected with probability 1 in large samples. This robustness property requires that we control for those lags of the data that are strong predictors of the outcome or impulse variables, but—crucially for applied work—the omission of lags with small-to-moderate predictive power does not threaten coverage. We argue that our result can be interpreted as a consequence of the double robustness of the LP estimator, which is analogous to the double robustness of modern partially linear regression estimators in the literature on debiased machine learning (e.g., see Newey, 1990; Ai and Chen, 2007; Chernozhukov, Chetverikov, Demirer, Duflo, Hansen, Newey, and Robins, 2018; Chernozhukov, Escanciano, Ichimura, Newey, and Robins, 2022).
@@
In stark contrast to LP, even small amounts of misspecification can cause conventional VAR confidence intervals for impulse responses to suffer from severe undercoverage. We first derive analytically the worst-case bias and coverage of VARs over all possible misspecification processes, subject to a constraint on the overall magnitude of the misspecification. From here a “no free lunch” result for VARs emerges: the worst-case bias and coverage distortion are small if, and only if, the asymptotic variance is close to that of LP. In general, the only way to guarantee robustness of conventional VAR inference is thus to include so many lags that the VAR estimator is asymptotically equivalent with LP. If instead the VAR confidence interval is much shorter (as is typically the case in applied practice), then VAR confidence intervals will severely undercover even for a misspecification term that: (i) is small in magnitude; (ii) has dynamic properties that cannot be ruled out ex ante based on economic theory; and (iii) is difficult to detect ex post with model specification tests. Instead of increasing the lag length, coverage can also be restored by using a larger bias-aware critical value (Armstrong and Kolesár, 2021), but we show that the resulting confidence intervals are so wide that one may as well report the LP interval.
@@
We demonstrate through simulations that our asymptotic results are useful to understand the finite-sample trade-off between LP and VAR confidence intervals. We consider a researcher that observes data generated from the Smets and Wouters (2007) model, and uses LPs and VARs to estimate the dynamic causal effects of cost-push or monetary shocks. If the lag length is selected by the Akaike Information Criterion (AIC), then VAR confidence intervals materially undercover—particularly at medium and long horizons—while LP throughout attains close to nominal coverage. Consistent with our theoretical results, increasing the estimation lag length ameliorates the VAR coverage, but at the cost of delivering confidence intervals as wide as those of LP.
Literature.
@@
Relative to the previously cited simulation studies of LPs and VARs, we here derive analytical results on the worst-case asymptotic properties of these two inference procedures that hold for a wide range of stationary, locally misspecified VAR models. The simulations in Li, Plagborg-Møller, and Wolf (2024) suggest a stark bias-variance trade-off between LP (low bias, high variance) and moderate-lag VAR estimators (moderate bias, low variance). The reason behind the theoretical superiority of LP proved in this paper is that, if the objective is to construct confidence intervals with robust coverage for a wide range of DGPs, then even a moderate amount of VAR bias cannot be tolerated, as it causes the VAR confidence interval to be poorly centered. A concern for correct confidence interval coverage thus effectively induces a large weight on bias in the researcher’s objective function, justifying the use of LP despite its higher variance.
@@
The robustness of LPs to misspecification discussed here—with stationary data and at fixed horizons—is conceptually and theoretically distinct from the robustness of LPs to the persistence in the data and the length of the impulse response horizon shown by Montiel Olea and Plagborg-Møller (2021). Nevertheless, it turns out that controlling for lags (“lag augmentation”) is key to all the robustness properties established in Montiel Olea and Plagborg-Møller (2021) and in the present paper.
@@
We also build upon previous research into misspecified VAR models, uncovering novel results about the robustness of LPs and the worst-case properties of VAR procedures. Braun and Mittnik (1993) derive expressions for the probability limits of VAR estimators under global MA misspecification; however, since bias always dominates variance asymptotically in their framework, they do not characterize the properties of LP and VAR inference procedures, which is the focus of our paper. Schorfheide (2005) characterizes the asymptotic mean squared errors of iterated and direct multi-step forecasts in a reduced-form VAR model with MA terms of order $T^{-1/2}$, and González-Casasús and Schorfheide (2024) use this framework to select hyperparameters for VAR forecasts. Müller and Stock (2011) construct Bayesian forecast intervals in a locally misspecified univariate AR model. Relative to these papers, we contribute by: (i) focusing on structural impulse responses rather than forecasting; (ii) allowing for more general rates of local misspecification, key to uncovering the double robustness of LP; and (iii) deriving simple analytical formulae for worst-case bias and coverage of VARs. As such, our results formalize concerns by applied practitioners about the lack of VAR robustness and sensitivity to lag length (Chari, Kehoe, and McGrattan, 2008; Nakamura and Steinsson, 2018; see also Kilian and Lütkepohl, 2017, Chapters 2.6.5 and 6.2).
@@
Whereas our paper deals with bias imparted by dynamic misspecification, the analysis does not capture other familiar sources of small-sample bias. In particular, our asymptotics abstract from the order-$T^{-1}$ biases arising from (i) persistence in the data (Pope, 1990; Kilian, 1998; Herbst and Johannsen, 2024) and (ii) the nonlinearity of the impulse response transformation of the VAR parameters (Jensen’s inequality).
@@
Section 2 defines the local-to-SVAR model and the LP and VAR estimators. Section 3 proves the robustness of LP and fragility of VAR confidence intervals. Section 4 derives analytically the worst-case bias and coverage of VARs, presents our “no free lunch” result, and shows that bias-aware VAR confidence intervals tend to be wider than the LP interval. Section 5 demonstrates the practical relevance of our results through simulations. Section 6 concludes. Replication codes are available online.
@@
2 Framework
We start out by defining the model and estimators.